# Artificial Precision Timing Array: bridging the decihertz gravitational-wave sensitivity gap with clock satellites

### v1

@@ IV.3 APTA sensitivity curves
Fundamentally, the sensitivity model involves five free parameters: $\rho_{th}$, $\delta t$, $N_{s}$, $T$ and $\sigma$. Following Ref. Moore et al. (2015a), we choose $\rho_{th}=3$ as a standard value for the rest of our study. We use $\delta t=0.01$ s, which is an order of magnitude below the minimum requirement for accessing 10 Hz GWs. We fiducially choose $N_{s}=6$ as a reasonable number of satellites.
@@
We assume that all the error in the timing residuals arises from the clock’s uncertainty, i.e. $\sigma=\xi\delta t$. Motivated by developments in clock technology (see Section IV.1), we choose three clock fractional uncertainties for our analysis: $\xi=10^{-17}$, the current record in 0.01 s of averaging Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017); $\xi=10^{-19}$, a realistic number for the next decade Ludlow et al. (2015); and $\xi=10^{-22}$, a high precision value included to establish an upper bound for $\xi$ in the foreseeable future.
@@
In Fig. 2, we display APTA sensitivity curves alongside example GW strains of merging intermediate-mass BBHs and early inspiraling stellar-mass and primordial BBHs. With current atomic clock technology, represented by $\xi=10^{-17}$ Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA could observe $10^{3}$ $\mathrm{M}_{\odot}$ and $10^{4}$ $\mathrm{M}_{\odot}$ BBH mergers at Gpc distances, as well as briefly see $100$ $\mathrm{M}_{\odot}$ binary inspirals. These inspirals are more clearly observed by APTA with the $\xi=10^{-19}$ near-future clock technology. Progress to $\xi=10^{-22}$ would allow the detection of $0.1$ $\mathrm{M}_{\odot}$ primordial-black-hole binaries. Alternatively, the current clock state of the art could see these sources using a lower $\delta t$ or if they are closer.
@@
Figure 2: Dimensionless characteristic GW strain as a function of frequency for different APTA clock precisions and binary black hole inspiral masses. The colored solid lines are APTA sensitivity curves for a total observation time $T=10$ s, pulse interval $\delta t=0.01$ s, number of satellites $N_{s}=6$, and threshold SNR $\rho_{th}=3$. The black lines represent binary black holes with mass ratio $q=1$ at a distance $D=1$ Gpc, with masses as indicated. The solid and dashed lines representing the binaries with total masses $M=10^{4}$ $\mathrm{M}_{\odot}$ and $M=10^{3}$ $\mathrm{M}_{\odot}$ end respectively at $f=1.3$ Hz and $f=13$ Hz, the approximate values for their mergers (see Eq. 2).
In Fig. 3, the sensitivity of APTA for the same three cases as in Fig. 2 is displayed together with the sensitivity curves of many current and future detectors, besides shaded areas corresponding to astrophysical GW sources Moore et al. ; Moore et al. (2015b). Sensitivity curve colors match those of Fig. 2, such that APTA-17 has $\xi=10^{-17}$, APTA-19 has $\xi=10^{-19}$, and APTA-22 has $\xi=10^{-22}$. With current clock technology Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA can surpass TianQin and LISA’s sensitivity in the dHz band and see some compact-binary inspirals. With near-future clock precision Ludlow et al. (2015), APTA may additionally access extreme-mass-ratio inspirals and GW bursts from Type Ia supernovae. Far future clocks with $\xi=10^{-22}$ can allow APTA to considerably outperform other proposed detectors DECIGO Kawamura et al. (2006) and BBO Phinney et al. (2004); Harry et al. (2006).
Figure 3: Dimensionless characteristic GW strain as a function of frequency for different astrophysical sources (shaded areas) and detectors (curves) Moore et al. ; Moore et al. (2015b). We display APTA for $[\delta t,T,\rho_{th},N_{s}]=[0.01\mathrm{s},10\mathrm{s},3,6]$. APTA-17 in purple has $\xi=10^{-17}$, APTA-19 in blue has $\xi=10^{-19}$, and APTA-22 in red has $\xi=10^{-22}$. The other detectors are Virgo Accadia et al. (2012), Advanced LIGO (aLIGO) Aasi et al. (2015), Einstein Telescope (ET) Punturo et al. (2010), Cosmic Explorer (CE) Reitze et al. (2019), TianQin Luo et al. (2016), the DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) Kawamura et al. (2006), Big Bang Observer (BBO) Phinney et al. (2004); Harry et al. (2006), and LISA Amaro-Seoane et al. (2017).
V Timing model Considerations
@@
Several additional effects must be taken into account in a realistic timing model. Although detailed calculation of these effects is beyond the scope of this study, we discuss some of them below.
@@
Shapiro time delay, the change in the travel time of light between two points due to the presence of massive bodies Shapiro (1964), from various solar system objects will introduce delays of $2GM/c^{3}$, e.g. $\sim 10^{-5}$ for the Sun, $\sim 10^{-8}$ s for Jupiter, and $\sim 10^{-13}$ s for the Moon. Since the main asteroid belt has an estimated total mass of $18\times 10^{-10}$ $\mathrm{M}_{\odot}$ Krasinsky et al. (2002), asteroids may also contribute with delays $\sim 10^{-14}$ s. These known delays may need to be included in the timing model, depending on where the satellites are placed. The gravity gradient noise of asteroids, on the other hand, only contributes marginally at much lower frequencies ($10^{-9}–10^{-5}$ Hz) Fedderke et al. (2022); Sullivan et al. (2022), and so should not affect APTA.
After setting the satellites in orbit, their positions should be tracked with sufficient precision to account for their motion and added time delays. Additionally, it will be important to track disruptions to the network (e.g. a satellite being captured by a planet or colliding with an asteroid). Disruptions to one or a few elements will not necessarily prevent GW detection since the signals detected by the other satellites remain correlated. If further analyses reveal that this is an important concern, the detector can be designed with a redundantly large number of satellites.
@@
Because of APTA’s small solar-system scale, some components of traditional PTA models, including parallax Deng and Finn (2011) as well as Einstein and Roemer delays (related to the use of a coordinate system centered in the solar system barycenter Hobbs et al. (2006)), are not as important. The solar wind Marsch (2006) contributes to the dispersion measure (DM), i.e. the integrated column density of free electrons between the observer and a light source. While accounting for the effects of the time-dependent DM on the ToAs is not straightforward Agazie et al. (2023b), this problem is easier for APTA than for PTAs, since the only contributions are from the solar system rather than the interstellar medium.
@@
To first approximation, clock synchronization should not play an important role in the functioning of APTA. Regular PTAs are of course not synchronized clocks, and have already been proven successful by relying on the long-term timing precision of the individual pulsars. Still, in general, significant time precision losses may happen in the process of time transfer (Caldwell et al., 2023, e.g.). Preserving, transferring, and utilizing the satellite clocks’ precision is also mission-critical for success. Therefore, the generation of the electromagnetic pulses and their locking to the clock should preserve the accuracy of the time reference. This can be done with state-of-the-art frequency comb technology Udem et al. (2002); Diddams et al. (2020). Additionally, the pulses’ phase information and intensity must not degrade sufficiently during their travel to compromise the recovery of the timing information. The multi-beam detection scheme at the central location must also preserve the mission-critical relative timing information. Each of these interwoven problems will require its own individual technical analysis as the required clock accuracy only sets the performance specifications of the downstream systems engineering.
@@ VI Conclusion
We have presented a proof-of-principle study into the use of an artificial solar-system-scale timing array to detect 0.1–10 Hz GWs, focusing on the required time reference precision. APTA would consist of $\sim 10$ satellites equipped with precision clocks orbiting and preferably communicating with a near-Earth base so the atmosphere is not in the path. The desired frequency band can be reached by targeting astrophysical signals at least 10 s long and ensuring that the beacons’ pulses are received at least every 0.1 s.
APTA’s sensitivity is fundamentally limited by $\sigma$, the error in the timing residuals. With the current state-of-the-art atomic clock technology Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA can achieve better dHz sensitivity than LISA and TianQin, allowing for detection of $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ black hole mergers and the early inspirals of some ground-based GW sources. With future clock technology available Ludlow et al. (2015), APTA could compete with and eventually surpass the sensitivity of other proposed detectors for the 0.1–10 Hz frequency range without the same challenges in high-precision interferometry they face. Then, it could detect most astrophysical sources in this band. These exciting prospects motivate the continued development of atomic clock technology and its application to GW detection.
While focusing on the accuracy of the APTA’s time reference, we have only discussed briefly the effects of additional noise sources, technical challenges, and technical developments required. This will require more detailed future work. Nevertheless, with conservative choices for measurement cadence $\delta t$, we have shown that high sensitivity can be still be achieved. Future studies may indicate that $\delta t$ can be lowered without straining APTA’s engineering requirements. With this and future papers laying the foundation, we propose APTA to observe uncharted territory of the gravitational-wave sky and gain new insight into high-energy astrophysics and cosmology.
@@ Acknowledgements
The authors thank Columbia University in the City of New York and Stanford University for their generous support. The Columbia Experimental Gravity group is grateful for the generous support of Columbia University. LMBA is grateful for the Columbia Undergraduate Scholars Program Summer Enhancement Fellowship and the Columbia Center for Career Education Summer Funding Program. AS is grateful for the support of the Stanford University Physics Department Fellowship and the National Science Foundation Graduate Research Fellowship.

### v2

@@ IV.2 APTA sensitivity curves
Fundamentally, the sensitivity model involves six free parameters: $\rho_{th}$, $N_{s}$, $\delta t$, $T$, $\sigma$, and $d_{s}$. We choose $\rho_{th}=10$ as a standard value for the rest of our study and fiducially pick $N_{s}=6$. To focus on the dependence on the clock uncertainty as previously discussed, we use $1/\delta t=10$ kHz.
@@
We assume that all the error in the timing residuals arises from the clock’s uncertainty, i.e. $\sigma=\xi_{\delta t}\delta t=\xi_{1\text{ s}}\sqrt{\delta t}$. Motivated by developments in clock technology (see Section IV.1), we choose three cases for the clock fractional uncertainty at 1 s of averaging: $10^{-18}$ representing present-day clocks, $10^{-20.5}$ representing near-future clocks, and $10^{-23}$ representing far-future clocks Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017).
The model in Section III concerns the sensitivity of PTAs to monochromatic GW sources, a proxy for slowly evolving binaries. We use $q=1$, $D=1$ Gpc BBHs as discussed in Section II as example target sources for APTA with total masses of $10^{4}$, $10^{3}$, $10^{2}$, and $10^{-1}$ $\mathrm{M}_{\odot}$. In the first 10 s after entering the APTA band, the heaviest (and thus fastest-evolving) binary’s GW frequency $f$ changes by $\sim 1.5\%$, so the monochromatic approximation holds. We initially use $T=10$ s to conservatively cover all four cases.
[…]
Figure 2: Dimensionless characteristic strain as a function of APTA radius $d_{s}$ for different GW frequencies. Other input parameters for APTA sensitivity curves are a total observation time $T=10$ s, pulse interval $\delta t=10^{-4}$ s, number of satellites $N_{s}=6$, and threshold SNR $\rho_{th}=10$.
@@
In Fig. 3, we display APTA sensitivity curves alongside example GW strains of merging intermediate-mass BBHs and early inspiraling stellar-mass and primordial BBHs. With current atomic clock technology, represented by $\xi_{1\text{ s}}=10^{-18}$ Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA could already observe $10^{3}$ $\mathrm{M}_{\odot}$ and $10^{4}$ $\mathrm{M}_{\odot}$ BBHs at or close to merger at Gpc distances, and, for $d_{s}=d_{\rm LISA}$, see a part of the evolution of 100 $\mathrm{M}_{\odot}$ binary inspirals before they enter the LIGO band. With progress in the next decade to values of $\xi_{1\text{ s}}$ around $10^{-20.5}$, the 0.1 $\mathrm{M}_{\odot}$ PBH binary would become observable by the larger APTA, and better clocks could make the detection of lighter and lighter PBHs possible.
@@
Figure 3: Dimensionless characteristic GW strain as a function of frequency for different APTA configurations and binary black hole total masses. The colored lines are APTA sensitivity curves with $T=10$ s, $\delta t=10^{-4}$ s, $N_{s}=6$, and $\rho_{th}=10$; solid ones have $d_{s}=d_{\rm LISA}$ while dashed ones have $d_{s}=d_{\rm LISA}/100$; yellow indicates $\xi_{1\text{ s}}=10^{-18}$, red, $\xi_{1\text{ s}}=10^{-20.5}$, and purple, $\xi_{1\text{ s}}=10^{-23}$. The black and gray lines represent binary black holes with mass ratio $q=1$ at a distance $D=1$ Gpc, with masses as indicated. The solid and dash-dotted black lines representing the binaries with total masses $10^{4} M\mathrm{M}_{\odot}$ and $10^{3} M\mathrm{M}_{\odot}$ end at their respective merger frequencies (see Eq. 2).
In Fig. 4, we showcase the sensitivity of the LISA-sized APTA alongside many future detectors in terms of the characteristic strain $h_{c}$ and the gravitational-wave energy density parameter per logarithmic frequency $\Omega_{\rm GW}(f)=2\pi^{2}f^{2}h_{c}(f)^{2}/3H_{0}^{2}$ Moore et al. (2015b), where $H_{0}=67.4$ km/s/Mpc is the Hubble constant Aghanim et al. (2020), related to the current rate of expansion of the universe. We focus on proposed detectors with the potential of observing astrophysical sources in the decihertz band. For this figure, $T$ is raised to $10^{4}$ s since precision timing can only be used to observe GWs with $f\gtrsim 1/T$. While, for some sources, this is inconsistent with our assumption of monochromatic waves, it is still a reasonable timescale for the observation of compact binary inspirals and other sources in APTA’s band, and a realistic APTA detector should certainly be modeled and constructed to follow GWs whose frequencies evolve in time.
With current clock technology Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA can achieve better sensitivity than Taiji, LISA, and TianQin for all frequencies displayed in Fig. 4 and outperform the atomic clock concept, ALIA, LGWA, DECIGO, and BBO in some frequency ranges. Progress toward lower clock uncertainties would allow APTA to surpass all of its existing and proposed peers for all frequencies.
In the right panel of Fig. 4, we also include the LIGO-Virgo-KAGRA prediction for the stochastic gravitational-wave background (SGWB) produced by BBHs and binary neutron stars (BNSs) Abbott et al. (2021b). Inspiraling compact binaries are known to create an SGWB spectrum $\Omega_{\rm SGWB}(f)\propto f^{2/3}$ Phinney (2001). We ignore deviations from this due to close-to-merger behavior and use the $2/3$ spectral index across all frequencies. Although we have not modeled APTA’s sensitivity to an SGWB, but solely to localized sources, the comparison between our sensitivity curves and the predicted SGWB signal remains informative because we only expect $\mathcal{O}(1)$ factors to appear following a more detailed analysis of a stochastic-background signal.
Figure 4: Dimensionless characteristic GW strain as a function of frequency for different detectors. We display APTA for $T,\delta t,\rho_{th},N_{s},d_{s}]=[10^{4}\text{ s},10^{-4}\text{ s},10,6,2.5\times 10^{6}\text{ km}]$ with different colors indicating different values of $\xi_{1\text{ s}}$ as discussed in the text. Other detectors, shown in black, are Taiji Ruan et al. (2020), TianQin Luo et al. (2016), LISA Amaro-Seoane et al. (2017), an atomic-clock-based technique Kolkowitz et al. (2016); Sedda et al. (2020), ALIA Bender (2004); Bender and Begelman (2005), LGWA Ajith et al. (2024), DECIGO Kawamura et al. (2006), and BBO Phinney et al. (2004); Harry et al. (2006). In the right-hand-side plot, we additionally display in blue the LIGO-Virgo-KAGRA prediction for the SGWB produced by BBHs and BNSs in our universe Abbott et al. (2021b).
V Towards a more realistic model of APTA
@@
This study’s main goal is to estimate the fundamental capabilities of detecting GWs with artificial “pulsar” timing. In a more realistic model of APTA, several additional effects must be taken into account. Although detailed calculation of these effects is beyond our present scope, we discuss some of them below.
@@
Shapiro time delay, the change in the travel time of light between two points due to the presence of massive bodies Shapiro (1964), from various solar system objects will introduce delays of order $2GM/c^{3}$, e.g. $\sim 10^{-5}$ s for the Sun, $\sim 10^{-8}$ s for Jupiter, and $\sim 10^{-13}$ s for the Moon. Since the main asteroid belt has an estimated total mass of $18\times 10^{-10}$ $\mathrm{M}_{\odot}$ Krasinsky et al. (2002), asteroids may also contribute with delays $\sim 10^{-14}$ s. These known delays will be included in the timing model. The gravity gradient noise of asteroids, on the other hand, only contributes marginally at much lower frequencies ($10^{-9}–10^{-5}$ Hz) Fedderke et al. (2022); Sullivan et al. (2022), so should not affect APTA.
In this study, we have utilized monochromatic binaries as example target sources for APTA. In reality, sources evolve in time. By not integrating the signal over the entire frequency band and not using any information about the signal’s time dependence, we underestimate the sensitivity of APTA. In future analyses, accounting for the evolution of signals in time will allow for the determination of how much better APTA’s sensitivity could be in an actual implementation.
In reality, APTA’s sensitivity curves in Figs. 3 and 4 should lose sensitivity at very low frequencies. Our simple models do not predict this because we have assumed all noise is clock-timing noise with a fixed uncertainty. A more realistic model would include error in the timing residual $\sigma$, cadence $1/\delta t$, and observing time $T$, which can vary from satellite to satellite and observation to observation Moore et al. (2015a), as well as various unresolved astrophysical sources at low frequencies Hils et al. (1990). This is a subject of future study.
@@
Because of APTA’s small solar-system scale, some components of traditional PTA models, including parallax Deng and Finn (2011) as well as Einstein and Roemer delays (related to the use of a coordinate system centered in the solar system barycenter Hobbs et al. (2006)), are not relevant. The solar wind Marsch (2006) contributes to the dispersion measure (DM), i.e. the integrated column density of free electrons between an observer and a light source. While accounting for the effects of the time-dependent DM on the ToAs is not straightforward Agazie et al. (2023b), this problem is easier for APTA than for PTAs, since contributions only come from the solar system and not from the interstellar medium.
After setting the satellites in orbit, their positions should be tracked with sufficient precision to account for their motion and added time delays. Additionally, it will be important to track disruptions to the network (e.g. a satellite being perturbed by a planet or colliding with an asteroid). Disruptions to one or a few elements will not necessarily prevent GW detection since the signals detected by the other satellites remain correlated. If further analyses reveal that this is an important concern, the detector can be designed with a redundantly large number of satellites.
@@
To first approximation, clock synchronization should not play an important role in the functioning of APTA. Regular PTAs are of course not synchronized clocks, and have already been proven successful by relying on the long-term timing precision of the individual pulsars. Still, in general, significant time precision losses may happen in the process of time transfer (see e.g. Caldwell et al. (2023)). Preserving, transferring, and utilizing the satellite clocks’ precision is also mission-critical for success. Therefore, the generation of the electromagnetic pulses and their locking to the clock should preserve the accuracy of the time reference. This can be done with state-of-the-art frequency comb technology Udem et al. (2002); Diddams et al. (2020). Additionally, the pulses’ phase information and intensity must not degrade sufficiently during their travel to compromise the recovery of the timing information. The multi-beam detection scheme at the central location must also preserve the mission-critical relative timing information. Each of these interwoven problems will require its own individual technical analysis as the required clock accuracy only sets the performance specifications of the downstream systems engineering.
@@ VI Conclusion
We have presented a proof-of-principle study into the use of an artificial precision timing array to detect 0.1–10 Hz GWs, focusing on the required time reference precision. APTA would consist of a few satellites equipped with precision clocks orbiting and communicating preferably with a space-based central station so the atmosphere is not in the path. The desired frequency band can be reached by targeting astrophysical signals at least 10 s long and ensuring that the beacons’ pulses are received at least every 0.1 s. APTA would not need to be larger than LISA and already performs significantly well with $1\%$ of LISA’s size.
Compared to traditional PTAs, APTA could enable better GW source localization, mainly because of the high precision in the knowledge of its satellites’ positions and lower clock uncertainty. Modern solar system ephemerides such as DE440 are advanced enough that they produce positional errors of less than 1 km for most celestial bodies across the solar system Park et al. (2021). This hints at great prospects for APTA compared to traditional PTAs, which have pulsar positional uncertainties $\gtrsim 100$ pc. The effect of this improved precision has been studied numerically for the pulsar timing case, where the angular uncertainty area decreased by almost 5 orders of magnitude as precision increased from 100 pc to 0.01 pc. In addition, when compared to typical PTA timing uncertainties of $\lesssim 100$ ns, APTA’s higher-precision clocks could considerably reduce the angular uncertainty area Kato and Takahashi (2023).
APTA’s sensitivity is fundamentally limited by $\sigma$, the error in the timing residuals. With the current state-of-the-art atomic clock technology Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA could achieve better dHz sensitivity than LISA, TianQin, and Taiji, allowing for the detection of $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ black hole mergers and the early inspirals of some ground-based GW sources. With future clock technology available Ludlow et al. (2015), APTA could compete with and eventually surpass the sensitivity of other proposed detectors for the 0.1–10 Hz frequency range without the same challenges in high-precision interferometry they face. Then, it could detect most astrophysical sources in this band. These exciting prospects motivate the continued development of precision clock technology and its application to GW detection.
While focusing on the accuracy of APTA’s time reference, we have only discussed briefly the effects of additional noise sources, technical challenges, and technical developments required. This will require more detailed future work. Nevertheless, with conservative choices for measurement cadence $\delta t$, we have shown that high sensitivity can be still be achieved, raising optimism for the APTA concept. With this and future papers laying the foundation, we propose APTA to observe uncharted territory of the gravitational-wave sky and gain precious new insight into high-energy astrophysics and cosmology.
@@ Acknowledgements
The authors thank Columbia University in the City of New York and Stanford University for their generous support. LMBA is grateful for the Columbia Undergraduate Scholars Program Summer Enhancement Fellowship, the Columbia Center for Career Education Summer Funding Program, and the support from Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES). AS is grateful for the support of the Stanford University Physics Department Fellowship, the National Science Foundation Graduate Research Fellowship, and a Giddings Fellowship to the Kavli Institute for Particle Astrophysics and Cosmology at Stanford. DV is grateful to The Scientific and Technological Research Council of Türkiye for their support through the grants 123C484 and 123C213; and to the European Commission through the MSCA COFUND Programme with the Grant Agreement No. 101081645. Especially, we would like to thank Peter Bender, Rainer Weiss, Stephen Taylor, Gianluca Giavitto, and Marek Kowalski for their ideas, encouragement, help, and suggestions.